A linear spring has a free length of 56.6 cm. When it is under a hanging load of 22949N, its total length is 1214.3 cm.  


  1. the spring constant k is  
  2. the hanging load that makes it 1749.7 cm long is  
  3. the energy stored in the spring when the extension is 1749.7 cm is  
1
Expert's answer
2020-11-20T07:12:56-0500

According to the Hook's law,

"F = k \Delta l"

where F is load, k - stiffness and "\Delta l = l_0 - l" - length difference.

  1. "\displaystyle k = \frac{F}{\Delta l} = \frac{22949}{(1214.3 - 56.6) \cdot 10^{-2} }= 1982.3 \; \;\frac{N}{m}"
  2. "F = 19 82.3 \cdot (1749.7-56.6) \cdot 10^{-2}= 33 562 \; N"
  3. "\displaystyle E = \frac{k (\Delta l)^2}{2} = \frac{1982.3 \cdot (1749.7-56.6)^2 \cdot 10^{-4}}{2} = 284\,122\; J"